OpenStudy (anonymous):

Simplifying radicals help? Simplify 2/√5 A. 2/√5 B. √10 C. √10/5 D. 2√5/5 Simplify -11√122 A. -44√7 B. -176√7 C. -27√7 D. 4√7 Simplify 17√17 - 9√17 A. 8 B. 8√17 C. 26√17 D. 8√0 I just want the explanation to solve these, no one can tell me! :(

jhonyy9 (jhonyy9):

so this mean rationalizing denominator yes ?

jhonyy9 (jhonyy9):

so for example 1/sqrt3 = sqrt3 /3 ok ?

jhonyy9 (jhonyy9):

for you can eliminate the radical from denominator you nee multiplie the numerator and the denominator by the radical from denominator because on this way will get on denominator this radical squared what mean that the radical is eliminated on denominator ok ? hope that is understandably

jhonyy9 (jhonyy9):

so this mean that for the first exercise is right answer D. ok ?

jhonyy9 (jhonyy9):

2 2 sqrt5 2sqrt5 2sqrt5 ----- = ------ * ------- = ------------ = ------- sqrt5 sqrt5 sqrt5 sqrt5 *sqrt5 5 hope that now you can understandig right easy ok ?

OpenStudy (anonymous):

So you multiply the 2 numbers on each side of it?

OpenStudy (mertsj):

\[\frac{2}{\sqrt{5}}\times \frac{\sqrt{5}}{\sqrt{5}}=\frac{2\sqrt{5}}{5}\]

OpenStudy (mertsj):

The second one, \[-11\sqrt{122}\] is already in simple form. I can only assume there is a typo.

OpenStudy (mertsj):

For the third one...if you want to add or subtract radicals, the radicands must be the same. If they are the same, simply add or subtract the coefficients so: \[17\sqrt{17}-9\sqrt{17}=8\sqrt{17}\]

OpenStudy (anonymous):

@mertsj & @jhonyy9 What about the second one? Its not a typo, but I am slowly catching on to this

OpenStudy (mertsj):

Is the second one: \[11\sqrt{121}\]